Corporation offers shares at the following par values:

            Common @ $3.00/par

            Preferred @ $6.50/par

1) Sylvester spends a total of $1725 for Common and Preferred shares. He has purchased $975 worth of Preferred. How many shares has he purchased of Common shares? How many shares of Preferred?

2). After his initial purchases Sylvester purchases 1/5 more of Common and 1/3 more of preferred. How much more did he spend on the purchases? What is the total of new shares owned?

3). Corporation offers a stock split on Common of 3:1. What is the par value of Sylvester’s Common stock after the split? What is the total value of shares owned by Sylvester?

4.) Sylvester decides to sell 15% of his Common shares and 20% of his preferred shares at their perspective share values. How much does he earn total on the sale?

5.) Following the sale of his shares Corporation offers a stock dividend on Common @$1.75/share and on Preferred @$2.19/share. How much is Sylvester’s total dividend?
Answer #1

1.
Common Shares purchased = (1725-975) / 3 = 250 shares
Preferred Shares purchased = 975 / 6.50 = 150 shares

2.
Common Shares Purchased = 250 x 1/5 = 50 shares @ $3 for $150
Preferred Shares Purchased = 150 x 1/3 = 50 shares @ $6.50 for $325
Total Amount spent = $475 i.e. $150 + 325

Total Common Shares owned = 250 + 50 = 300 shares
Total Preferred Shares owned = 150 + 50 = 200 shares
Total Shares = 500 shares

3.
Par value of common stock after stock split = $3 x 1/3 = $1 per share
Total Value of shares owned = $1725 + 475 = $2200

4.
It is not clear what does earn here refer to, profit or sales proceeds
Sales proceeds from sale = 300 x 3 x 15% x $1 + 200 x 20% x $6.50 = $395
Profit will be 0, since it is assumed shares are sold at par value

5.
Common Stock Dividend = 300 x 3 x 85% x $1.75 = $1338.75
Preferred Stock Dividend = 200 x 80% x $2.19 = $350.40
Total Dividend = $1689.15

If you have any query, kindly comment with your query and please mark thumbs up.